You analyze the voltages across different branches of a simple circuit (using Kirchoffs Laws), and you find the following equations relating the voltages across different resistors 4V 1+4V 2 90 6.1 V 2+7 v 3 72 where vi represents current through wire 1 2 through wire 2, etc. All voltages are in V (Volts). Using matrix methods, determine the magnitude of vi1 in Volts Answer Check
Answer #1

Solution by Inverse Matrix Method 3 90 5-1 100 36 39 35 296 148 74 35 -35 15 296 148 74 -61 61-5 592 296 148 -15 39 35 296 148 74 -15 5039 148 -1709 148 3283 296 X A-1x35-35 15 100 296 148 74 - 61 61 592 296 148v_1 = 5039/148 = 34.047 V

v_2=-1709/148 V

v_3=3283/296 V